solve: if 3 coffee and 2 donuts = 8.75 and 1 coffee and 2 donuts = 4.25 what is the cost of 1 coffee

Found 2 solutions by Boreal, ikleyn:
Answer by Boreal(15235)

3C+2D=8.75
1C+2D=4.25
-1C-2D=-4.25
2C=4.50
C=$2.25
3C+2D=6.75+2D=8.75
D=$1
Check in second
2.25+2=$4.25 checks
Coffee is $2.25
Doughnut is $1

Answer by ikleyn(52798)

.
The nearest consequence from the condition is that 2 coffee cost 8.75 - 4.25 = 4.50.

Can you complete from this point ?